# VCR-GauS: View Consistent Depth-Normal Regularizer for Gaussian Surface Reconstruction

### v1

@@ Abstract
Although 3D Gaussian Splatting has been widely studied because of its realistic and efficient novel-view synthesis, it is still challenging to extract a high-quality surface from the point-based representation. Previous works improve the surface by incorporating geometric priors from the off-the-shelf normal estimator. However, there are two main limitations: 1) Supervising normal rendered from 3D Gaussians updates only the rotation parameter while neglecting other geometric parameters; 2) The inconsistency of predicted normal maps across multiple views may lead to severe reconstruction artifacts. In this paper, we propose a Depth-Normal regularizer that directly couples normal with other geometric parameters, leading to full updates of the geometric parameters from normal regularization. We further propose a confidence term to mitigate inconsistencies of normal predictions across multiple views. Moreover, we also introduce a densification and splitting strategy to regularize the size and distribution of 3D Gaussians for more accurate surface modeling. Compared with Gaussian-based baselines, experiments show that our approach obtains better reconstruction quality and maintains competitive appearance quality at faster training speed and 100+ FPS rendering. Our code will be made open-source upon paper acceptance.
@@ 1 Introduction
Multi-view stereo (MVS) is a long-standing problem that aims to create 3D surfaces of an object or scene captured from multiple viewpoints [9; 5; 24; 40]. This technique has applications in robotics, graphics, virtual reality, etc. Recently, rendering methods [48; 58; 25] have enhanced the quality of reconstructions. These approaches, often based on implicit neural representations, require extensive training time. For instance, Neuralangelo [25] uses hash encoding [32] for creating high-fidelity surfaces but requires 128 GPU hours for a single scene. On the other hand, the novel 3D Gaussian Spatting method [21] employs 3D Gaussians to render complex scenes photorealistically in real-time, offering a more efficient alternative. Consequently, many recent works attempted the utilization of Gaussian Splatting for surface reconstruction [8; 15; 44; 18]. Although they achieve success in object-level reconstruction, it is still challenging to extract a high-quality surface for large scenes. Previous works [58] improve the surface for scene-level reconstruction by incorporating geometric priors from the off-the-shelf normal estimator. However, there are two main limitations for Gaussian-based reconstruction: 1) Supervising normal rendered from Gaussian Splatting by its nature can only update the rotation parameters, hence failing to update other geometric parameters; 2) The predicted normal maps are inconsistent across multiple views, which may lead to severe reconstruction artifacts.
@@
In this paper, we introduce a novel view-consistent Depth-Normal (D-Normal) regularizer to alleviate the above-mentioned two limitations. For the first limitation, we notice that the supervision of the Gaussian normals can only update its rotations without directly affecting its positions. Consequently, the supervision of Gaussian normals is not as effective as NeuS-based methods [48; 58; 25] whose normal is the gradient of the signed distance function (SDF) that is directly related to the position in 3D space. To solve this issue, we are inspired by the depth and normal estimation [1; 55] to introduce a D-Normal formulation, where the normal is derived from the gradient of rendered depth instead of directly blended from 3D Gaussians. Unlike existing works that obtain depth from the center position of 3D Gaussians, we compute the depth as the intersection of the ray and the compressed Gaussians. Specifically, we first make the Gaussians suitable for 3D reconstruction by applying a scale regularization similar to NeuSG [8] to compress the 3D Gaussian ellipsoids into a plane. Subsequently, the computation of the depth can be simplified to the intersection between a ray and a plane. As a result, our novel parametrization of the depth allows effective full supervision of the Gaussian geometric parameters by any data-driven monocular normal estimator.
@@ 2 Related Work
Novel View Synthesis. The pursuit of novel view synthesis began with Soft3D [33], which integrated deep learning and volumetric ray-marching to form a continuous, differentiable density field for geometry representation [17; 41]. While effective, this approach was computationally expensive. Neural Radiance Fields (NeRF) [31] improved render quality with importance sampling and positional encoding, but the deep neural networks slowed down processing. Subsequent methods aimed to optimize both quality and speed. Techniques like position encoding and band-limited coordinate networks are combined with neural radiance fields for pre-filtered scene representation [2; 3; 27]. Innovations to speed up rendering included leveraging spatial data structures and adjusting MLP size [6; 11; 14; 16; 34; 43]. Notable examples are InstantNGP [32], which uses a hash grid and a reduced MLP for faster computation, and Plenoxels [11], which employs a sparse voxel grid to eliminate neural networks entirely. Both use Spherical Harmonics to enhance rendering. Despite these advancements, challenges remain in representing empty space and maintaining image quality with structured grids and extensive sampling. Recently, 3D Gaussian Splatting (3DGS) [21] has addressed these issues with unstructured and GPU-optimized splatting, achieving faster and higher-quality rendering without neural components. In this work, we utilize the advantage of Gaussian Splatting to perform surface reconstruction and incorporate normal priors to guide the reconstruction, especially for large indoor and outdoor scenes.
@@
Multi-View Surface Reconstruction. Surface reconstruction is key in 3D vision. Traditional MVS methods [4; 9; 5; 24; 37; 40; 39] use feature matching for depth [4; 37] or voxel-based shapes [9; 5; 24; 40; 45]. Depth-based methods combine depth maps into point clouds, while volumetric methods estimate occupancy and color in voxel grids [9; 5; 28]. However, the finite resolution of voxel grids limits precision. Learning-based MVS modifies traditional steps such as feature matching [30; 47; 59], depth integration [36], or depth inference from images [19; 50; 60; 51; 57]. Further advancements [48; 52] integrated implicit surfaces with volume rendering, achieving detailed surface reconstructions from RGB images. These methods have been extended to large-scale reconstructions via additional regularization [58; 25]. Despite these impressive developments, efficient large-scale scene reconstruction remains a challenge. For example, Neuralangelo [25] requires 128 GPU hours for reconstructing a single scene from the Tanks and Temples Dataset [23]. To accelerate the reconstruction process, some works [15; 18] introduce the 3D Gaussian splitting technique. However, these works still fail in large-scale reconstructions. In this work, we focus on introducing normal regularization for large-scale reconstructions.
@@
3D Gaussian Splatting. Since 3DGS [21] was introduced, it has been rapidly extended to surface reconstruction. We highlight the distinctions between our method and concurrent works SuGaR [15], 2DGS [18], NeuSG [8], and DN-Splatter [46]. In contrast to SuGaR and 2DGS with unsatisfactory performance on large-scale scenes, our method focuses on introducing normal regularization to improve large-scale reconstructions. 2DGS obtaining 2D Gaussian primitives by setting the last entry of scaling factors to zero which is hard to optimize by original Gaussian Splatting technique as noted in  [64; 18], while our method utilizes scale regularization to flatten 3D Gaussians which are easier to optimize. NeuSG utilizes both 3D Gaussian splitting and neural implicit rendering jointly and extracts the surface from an SDF network, while our approach is faster and conceptually simpler by leveraging only Gaussian splatting for surface approximation. Although normal prior is also used for indoor scenes, DN-Splatter may show severe reconstruction artifacts due to their normal supervision can only update the rotation parameters and normal maps inconsistencies across multiple views. Moreover, we do not use the ground truth depth for supervision utilized by DN-Splatter. In comparison, our work is designed to solve both limitations.
@@ 3.1 Preliminaries: 3D Gaussian Splatting
A 3D Gaussian is defined by its position $\mathbf{p}$, quaternion $\mathbf{r}$, scaling factor $\mathbf{s}$, opacity $\alpha\in\mathbb{R}$, and color represented with spherical harmonics coefficients $\mathbf{H}\in\mathbb{R}^{k}$. For a given pixel, the combined color and opacity from multiple Gaussians are weighted by Eq. 1. The color blending for overlapping points is:
@@
where $\mathbf{c}_{i}$ and $\alpha_{i}$ denote the color and density of a point, respectively.
@@ 3.3 View-Consistent D-Normal Regularization
$\mathcal{L}_{\text{n}}=\|\hat{\textbf{N}}-\textbf{N}\|_{1}+(1-\hat{\textbf{N}}\cdot\textbf{N}),$ (8)
where the rendered normal map $\hat{\mathbf{N}}$ applies weighted-sum normal, different from color rendering using alpha blending. However, normal regularization alone is insufficient for surface reconstruction as compared with NeuS-based methods. As illustrated in Fig. 2, NeuS-based methods compute normals as the gradients of the SDF function calculated from input position p, and therefore the regularization of normal can affect the position directly. However, the normal is only directly related to the rotation of the Gaussian during Gaussian splitting, which means directly supervising the rendered normals cannot update the positions as shown in Fig. 2. To solve this problem and inspired by normal and depth estimation [1; 55], we propose a new depth-normal formulation. First, we render the depth map by weighted summing the depths:
@@
$\bar{\textbf{N}}_{d}(\textbf{n},\textbf{p})=\frac{\nabla_{v}\mathbf{d}(\textbf{n},\textbf{p})\times\nabla_{h}\mathbf{d}(\textbf{n},\textbf{p})}{|\nabla_{v}\mathbf{p}(\textbf{n},\textbf{p})\times\nabla_{h}\mathbf{d}(\textbf{n},\textbf{p})|}.$ (10)
@@
$w=\exp\{(\bar{\textbf{N}}_{d}\cdot\textbf{N}-1)/\gamma\},$ (12)
@@ 4 Experiments
4.1 Comparison
Method F1-score Time
Implicit NeuS 65.12 $>$10h
MonoSDF 81.64
Explicit 3DGS 50.79 $\leq$1h
SuGar 63.20
2DGS 64.36
Ours 78.17
@@ 4.2 Ablation Studies
We verify the effectiveness of different design choices on reconstruction quality, including regularization terms, intersection depth, and densification on the TNT dataset [23] and report the F1-score. We first examine the effect of our view-consistent D-Normal regularization. Our full model (Tab. 4 E) provides the best performance (0.40 F1-score). The performance drops 0.10 F1-score from 0.4 to 0.3 without the D-Normal regularizer (Tab. 4 A) while keeping rendered normal regularization. It proves that it is insufficent to supervise only the normal maps rendered from Gaussian Splatting. Furthermore, the result drops by 0.04 F1-score without the confidence (Tab. 4 B) and with the D-Normal regularizer. It demonstrates that confidence can mitigate the problem of inconsistency of the predicted normal maps. From Fig. 5, we can observe that disabling the confidence leads to an unsmooth surface. Both of these validate the effectiveness of the view-consistent D-Normal regularization. Additionally, the absence of intersection depth (Tab. 4 C) results in poor performance. Lastly, the performance increases from 0.33 F1-score to 0.40 with our densification and split (Tab. 4 D), proving small Gaussians represent surfaces better than large Gaussians.
@@ A.1 Semantic Surface Trimming
Figure 6: Qualitative ablation for the semantic trimming. The left is disabling the semantic trimming and the right is enabling the strategy. We can see that the proposed semantic trimming can prune the unwanted regions, like the sky in the left figure.
@@
A.2 Implementation Details

### v2

@@ Abstract
Although 3D Gaussian Splatting has been widely studied because of its realistic and efficient novel-view synthesis, it is still challenging to extract a high-quality surface from the point-based representation. Previous works improve the surface by incorporating geometric priors from the off-the-shelf normal estimator. However, there are two main limitations: 1) Supervising normals rendered from 3D Gaussians effectively updates the rotation parameter but is less effective for other geometric parameters; 2) The inconsistency of predicted normal maps across multiple views may lead to severe reconstruction artifacts. In this paper, we propose a Depth-Normal regularizer that directly couples normal with other geometric parameters, leading to full updates of the geometric parameters from normal regularization. We further propose a confidence term to mitigate inconsistencies of normal predictions across multiple views. Moreover, we also introduce a densification and splitting strategy to regularize the size and distribution of 3D Gaussians for more accurate surface modeling. Compared with Gaussian-based baselines, experiments show that our approach obtains better reconstruction quality and maintains competitive appearance quality at faster training speed and 100+ FPS rendering.
@@ 1 Introduction
Figure 2: Illustration of rendered normal supervision and the D-Normal regularizer. (a) As a result of the back-propagation through alpha-blending via Eq. 1, rendered normal supervision $\mathcal{L}_{\text{n}}$ moves Gaussians closer to ($\textbf{P}_{1}$) or away from ($\textbf{P}_{2}$) the intersecting ray. When the normal of a Gaussian is closer to the GT surface normal, the supervision pushes this Gaussian ($\textbf{P}_{1}$) towards the ray to increase its weight in the rendering equation, and vice-versa ($\textbf{P}_{2}$). (b) Such movement of Gaussians stops when the rendered normal loss $\mathcal{L}_{\text{n}}$ is equal to zero. In either case ((a) or (b)), the rendered normal loss cannot move Gaussian towards the surface. In contrast, (c) the D-Normal regularizer $\mathcal{L}_{\text{dn}}$ can move Gaussians towards or away from GT surface. $\textbf{P}_{1}$ and $\textbf{P}_{2}$ are the 3D positions corresponding to the mean depth of two neighboring pixels (rays) via Eq. 10. The D-Normal $\bar{\textbf{N}}_{d}$ is derived from $\textbf{P}_{1}$ and $\textbf{P}_{2}$ in Eq. 11. $\mathcal{L}_{\text{dn}}$ encourages $\bar{\textbf{N}}_{d}$ to align with the ground truth normal N, resulting in Gaussians moving towards or away from the surface.
@@
Multi-view stereo (MVS) is a long-standing problem that aims to create 3D surfaces of an object or scene captured from multiple viewpoints [9, 5, 25, 41]. This technique has applications in robotics, graphics, virtual reality, etc. Recently, rendering methods [49, 59, 26, 16] have enhanced the quality of reconstructions. These approaches which are often based on implicit neural representations require extensive training time. For instance, Neuralangelo [26] uses hash encoding [33] for creating high-fidelity surfaces but requires 128 GPU hours for a single scene. On the other hand, the novel 3D Gaussian Spatting method [22] employs 3D Gaussians to render complex scenes photorealistically in real-time, offering a more efficient alternative. Consequently, many recent works attempted the utilization of Gaussian Splatting for surface reconstruction [8, 15, 45, 19]. Although they achieve success in object-level reconstruction, it is still challenging to extract a high-quality surface for large scenes. Previous works [59] improve the surface for scene-level reconstruction by incorporating geometric priors from the off-the-shelf normal estimator. However, there are two main limitations for Gaussian-based reconstruction: 1) Supervising normals rendered from 3D Gaussians effectively updates the rotation parameter but is less effective for other geometric parameters; 2) The predicted normal maps are inconsistent across multiple views, which may lead to severe reconstruction artifacts.
@@
In this paper, we introduce a novel view-consistent Depth-Normal (D-Normal) regularizer to alleviate the above-mentioned limitations. As illustrated in Fig. 2, we notice that the supervision of the Gaussian normals can effectively update its rotations but is less effective for affecting its positions. Consequently, the supervision of Gaussian normals is not as effective as NeuS-based methods [49, 59, 26] whose normal is the gradient of the signed distance function (SDF) that is directly related to the position in 3D space. To solve this issue, we are inspired by the depth and normal estimation [1, 56] to introduce a D-Normal formulation, where the normal is derived from the gradient of rendered depth instead of directly blended from 3D Gaussians. Unlike existing works that obtain depth from the center position of 3D Gaussians, we compute the depth as the intersection of the ray and the compressed Gaussians. Specifically, we first make the Gaussians suitable for 3D reconstruction by applying a scale regularization similar to NeuSG [8] to compress the 3D Gaussian ellipsoids into a plane. Subsequently, the computation of the depth can be simplified to the intersection between a ray and a plane. As a result, our novel parametrization of the depth allows effective full supervision of the Gaussian geometric parameters by any data-driven monocular normal estimator.
@@ 2 Related Work
Novel View Synthesis. The pursuit of novel view synthesis began with Soft3D [34], which integrated deep learning and volumetric ray-marching to form a continuous, differentiable density field for geometry representation [18, 42]. While effective, this approach was computationally expensive. Neural Radiance Fields (NeRF) [32] improved render quality with importance sampling and positional encoding, but the deep neural networks slowed down processing. Subsequent methods aimed to optimize both quality and speed. Techniques like position encoding and band-limited coordinate networks are combined with neural radiance fields for pre-filtered scene representation [2, 3, 28]. Innovations to speed up rendering included leveraging spatial data structures and adjusting MLP size [6, 11, 14, 17, 35, 44]. Notable examples are InstantNGP [33], which uses a hash grid and a reduced MLP for faster computation, and Plenoxels [11], which employs a sparse voxel grid to eliminate neural networks entirely. Both use Spherical Harmonics to enhance rendering. Despite these advancements, challenges remain in representing empty space and maintaining image quality with structured grids and extensive sampling. Recently, 3D Gaussian Splatting (3DGS) [22] has addressed these issues with unstructured and GPU-optimized splatting, achieving faster and higher-quality rendering without neural components. In this work, we utilize the advantage of Gaussian Splatting to perform surface reconstruction and incorporate normal priors to guide the reconstruction, especially for large indoor and outdoor scenes.
@@
Multi-View Surface Reconstruction. Surface reconstruction is key in 3D vision. Traditional MVS methods [4, 9, 5, 25, 38, 41, 40] use feature matching for depth [4, 38] or voxel-based shapes [9, 5, 25, 41, 46]. Depth-based methods combine depth maps into point clouds, while volumetric methods estimate occupancy and color in voxel grids [9, 5, 29]. However, the finite resolution of voxel grids limits precision. Learning-based MVS modifies traditional steps such as feature matching [31, 48, 60], depth integration [37], or depth inference from images [20, 51, 61, 52, 58]. Further advancements [49, 53] integrated implicit surfaces with volume rendering, achieving detailed surface reconstructions from RGB images. These methods have been extended to large-scale reconstructions via additional regularization [59, 26]. Despite these impressive developments, efficient large-scale scene reconstruction remains a challenge. For example, Neuralangelo [26] requires 128 GPU hours for reconstructing a single scene from the Tanks and Temples Dataset [24]. To accelerate the reconstruction process, some works [15, 19] introduce the 3D Gaussian splitting technique. However, these works still fail in large-scale reconstructions. In this work, we focus on introducing normal regularization for large-scale reconstructions.
@@
3D Gaussian Splatting. Since 3DGS [22] was introduced, it has been rapidly extended to surface reconstruction. We highlight the distinctions between our method and concurrent works SuGaR [15], 2DGS [19], NeuSG [8], and DN-Splatter [47]. In contrast to SuGaR and 2DGS with unsatisfactory performance on large-scale scenes, our method focuses on introducing normal regularization to improve large-scale reconstructions. 2DGS obtaining 2D Gaussian primitives by setting the last entry of scaling factors to zero which is hard to optimize by original Gaussian Splatting technique as noted in  [65, 19], while our method utilizes scale regularization to flatten 3D Gaussians which are easier to optimize. NeuSG utilizes both 3D Gaussian splitting and neural implicit rendering jointly and extracts the surface from an SDF network, while our approach is faster and conceptually simpler by leveraging only Gaussian splatting for surface approximation. Although normal prior is also used for indoor scenes, DN-Splatter may show severe reconstruction artifacts due to their normal supervision can only update the rotation parameters and normal maps inconsistencies across multiple views. Moreover, we do not use the ground truth depth for supervision utilized by DN-Splatter. In comparison, our work is designed to solve both limitations.
@@ 3.1 Preliminaries: 3D Gaussian Splatting
A 3D Gaussian is defined by its position $\mathbf{p}$, quaternion $\mathbf{r}$, scaling factor $\mathbf{s}$, opacity $o\in\mathbb{R}$, and color represented with spherical harmonics coefficients $\mathbf{H}\in\mathbb{R}^{k}$. For a given pixel, the combined color and opacity from multiple Gaussians are weighted by Eq. 1. The color blending for overlapping points is:
@@
where $\mathbf{c}_{i}$ and $\alpha_{i}=o_{i}G(\textbf{x}_{i})$ denote the color and density of a point, respectively.
@@ 3.3 View-Consistent D-Normal Regularization
$\mathcal{L}_{\text{n}}$ $=\|\hat{\mathbf{N}}-\mathbf{N}\|_{1}+(1-\hat{\mathbf{N}}\cdot\mathbf{N}),$ (8)
where $\text{where }\hat{\mathbf{N}}$ $={\sum_{i\in M}\mathbf{n}_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})}/{\sum_{i\in M}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})}.$ (9)
However, normal regularization alone is insufficient for surface reconstruction as compared with NeuS-based methods. There are two main reasons for this: 1) Updating the position of a Gaussian using $G(\textbf{x})$ only moves it closer to or farther from the intersecting ray, as shown in Fig. 2 (a) (the mathematical proof is provided in A.2 of the supplemental material); 2) NeuS-based methods calculate normals as gradients of the SDF function from the input position p and therefore normal regularization effectively influences position updates. However, since the normal is only related to the rotation of the Gaussian in 3DGS, supervising the rendered normals does not efficiently update positions as shown in Fig. 3. To solve this problem and inspired by normal and depth estimation [1, 56], we propose a new depth-normal formulation. First, we render the depth map by weighted summing the depths:
@@
$\bar{\textbf{N}}_{d}(\textbf{n},\textbf{p})=\frac{\nabla_{v}\mathbf{d}(\textbf{n},\textbf{p})\times\nabla_{h}\mathbf{d}(\textbf{n},\textbf{p})}{|\nabla_{v}\mathbf{d}(\textbf{n},\textbf{p})\times\nabla_{h}\mathbf{d}(\textbf{n},\textbf{p})|}.$ (11)
@@
$w=\exp\{(\hat{\textbf{N}}_{d}\cdot\textbf{N}-1)/\gamma\},$ (13)
@@ 4 Experiments
4.1 Comparison
Method F1-score Time
Implicit NeuS 65.12 $>$10h
MonoSDF 81.64
Explicit 3DGS 50.79 $\leq$1h
SuGar 63.20
2DGS 64.36
Ours 78.17
@@ 4.1 Comparison
Figure 6: Qualitative ablation for the D-Normal regularizer. We plot the positions of Gaussian centers from the optimized 3D scenes. The left disables the D-Normal regularizer with only rendered normal supervision, and the right enables both D-Normal and rendered normal supervision. Compared to w/o D-Normal that produces many noisy Gaussians floating off the surface, our proposed D-Normal regularizer effectively pushes the 3D Gaussians towards the surface and thus providing much cleaner reconstruction.
@@ 4.2 Ablation Studies
We verify the effectiveness of different design choices on reconstruction quality, including regularization terms, intersection depth, and densification on the TNT dataset [24] and report the F1-score. We first examine the effect of our view-consistent D-Normal regularization. Our full model (Tab. 4 E) provides the best performance (0.40 F1-score). The performance drops 0.10 F1-score from 0.4 to 0.3 without the D-Normal regularizer (Tab. 4 A) while keeping rendered normal regularization. It proves that it is insufficent to supervise only the normal maps rendered from Gaussian Splatting. The visualization in Fig. 6 demonstrates that our d-normal regularization can effectively push the 3D Gaussians towards the surface. Furthermore, the result drops by 0.04 F1-score without the confidence (Tab. 4 B) and with the D-Normal regularizer. It demonstrates that confidence can mitigate the problem of inconsistency of the predicted normal maps. From Fig. 7, we can observe that disabling the confidence leads to an unsmooth surface. Both of these validate the effectiveness of the view-consistent D-Normal regularization. Additionally, the absence of intersection depth (Tab. 4 C) results in poor performance. Lastly, the performance increases from 0.33 F1-score to 0.40 with our densification and split (Tab. 4 D), proving small Gaussians represent surfaces better than large Gaussians.
@@ 5 Conclusion
Acknowledgement. This research / project is supported by the National Research Foundation (NRF) Singapore, under its NRF-Investigatorship Programme (Award ID. NRF-NRFI09-0008).
@@ A.1 Semantic Surface Trimming
Figure 8: Qualitative ablation for the semantic trimming. The left is disabling the semantic trimming and the right is enabling the strategy. We can see that the proposed semantic trimming can prune the unwanted regions, e.g. the sky in the left figure.
@@
A.2 Proof on Our D-Normal Regularizer
Fig. 2 is to illustrate the optimization of positions of Gaussians under normal and d-normal supervisions. As we mentioned in Sec. 1, in contrast to supervision on rendered normal maps which only updates Gaussian rotations, our D-Normal regularizer can also effectively update the positions of the Gaussians. We show the mathematical proof below.
Proposition 1.1: Supervision on rendered normal cannot effectively affect the positions of Gaussians.
Proposition 1.2: Supervision on our D-Normal regularizer can effectively affect the positions of Gaussians.
Proof: Without a loss of generality, we omit the summation over multiple views in our following derivations for brevity. Based on the loss $\mathcal{L_{\text{n}}}$ on rendered normal (cf. Eq. 8), the gradient of $\mathcal{L_{\text{n}}}$ with respect to position $\mathbf{p}$:
$\frac{\partial\mathcal{L_{\text{n}}}}{\partial\mathbf{p}_{i}}$ $=\frac{\mathcal{L_{\text{n}}}}{\partial\hat{\mathbf{N}}}\cdot\frac{\partial\hat{\mathbf{N}}}{\partial\mathbf{p}_{i}},$ (16)
$\frac{\partial\hat{\mathbf{N}}}{\partial\mathbf{p}_{i}}$ $=\frac{\partial\hat{\mathbf{N}}}{\partial\alpha_{i}}\cdot\frac{\partial\alpha_{i}}{\partial\mathbf{p}_{i}}+\frac{\partial\hat{\mathbf{N}}}{\partial\mathbf{n}_{i}}\cdot\frac{\partial\mathbf{n}_{i}}{\partial\mathbf{p}_{i}}$
$=\frac{\partial\hat{\mathbf{N}}}{\partial\alpha_{i}}\cdot\frac{\partial\alpha_{i}}{\partial G(\mathbf{x})}\cdot\frac{\partial G(\mathbf{x})}{\partial\mathbf{p}_{i}}$
$=\frac{\partial\hat{\mathbf{N}}}{\partial\alpha_{i}}\cdot\frac{\partial\alpha_{i}}{\partial G(\mathbf{x})}\cdot[-G(\mathbf{x})\cdot(\mathbf{R}\mathbf{S}\mathbf{S}^{\top}\mathbf{R}^{\top})^{-1}\cdot(\mathbf{x}-\mathbf{p}_{i})]$
$\approx\frac{\partial\hat{\mathbf{N}}}{\partial\alpha_{i}}\cdot\frac{\partial\alpha_{i}}{\partial G(\mathbf{x})}\cdot[-G(\mathbf{x})\cdot(\mathbf{x}-\mathbf{p}_{i})]$ (17)
Putting Eq. 17 into Eq. 16, we get:
$\frac{\partial\mathcal{L_{\text{n}}}}{\partial\mathbf{p}_{i}}$ $\approx\frac{\mathcal{L_{\text{n}}}}{\partial\hat{\mathbf{N}}}\cdot\frac{\partial\hat{\mathbf{N}}}{\partial\alpha_{i}}\cdot\frac{\partial\alpha_{i}}{\partial G(\mathbf{x})}\cdot[-G(\mathbf{x})\cdot(\mathbf{x}-\mathbf{p}_{i})]$
$=\beta\cdot\frac{\partial\alpha_{i}}{\partial G(\mathbf{x})}\cdot[-G(\mathbf{x})\cdot(\mathbf{x}-\mathbf{p}_{i})]$
$\propto(\mathbf{x}-\mathbf{p}_{i}),\text{where }\beta=\frac{\mathcal{L_{\text{n}}}}{\partial\hat{\mathbf{N}}}\cdot\frac{\partial\hat{\mathbf{N}}}{\partial\alpha_{i}}\text{is a scalar}.$ (18)
Based on the D-Normal regularization $\mathcal{L_{\text{dn}}}$ (cf. Eq. 12), the gradient of $\mathcal{L}_{\text{dn}}$ with respect to position $\mathbf{p}$:
$\frac{\partial\mathcal{L}_{\text{dn}}}{\partial\textbf{p}_{i}}$ $=\frac{\partial\mathcal{L}_{\text{dn}}}{\partial\bar{\textbf{N}}_{d}}\cdot\frac{\partial\bar{\textbf{N}}_{d}}{\partial\hat{D}}\cdot\frac{\partial\hat{D}}{\partial\textbf{p}_{i}},$
$\frac{\partial\hat{D}}{\partial\textbf{p}_{i}}$ $=\frac{\partial\hat{D}}{\partial\alpha_{i}}\cdot\frac{\partial\alpha_{i}}{\partial\textbf{p}_{i}}+\frac{\partial\hat{D}}{\partial d_{i}}\cdot\frac{\partial d_{i}}{\partial\textbf{p}_{i}}$
$=\frac{\partial\hat{D}}{\partial\alpha_{i}}\cdot\frac{\partial\alpha_{i}}{\partial G(\textbf{x})}\cdot\frac{\partial G(\textbf{x})}{\partial\textbf{p}_{i}}+\frac{\partial\hat{D}}{\partial d_{i}}\cdot r_{z}\cdot\frac{\mathbf{n}}{\mathbf{n}\cdot\mathbf{r}}.$ (19)
We can deduce the following from Eq. 16 and Eq. 19:
Case 1: From Eq. 16, we can see that the gradient-update $\frac{\partial\mathcal{L_{\text{n}}}}{\partial\mathbf{p}_{i}}$ of position is independent of the normal $\mathbf{n}$. Consequently, the supervision on rendered normal cannot effectively affect the Gaussian position $\mathbf{p}$.
Case 2: From Eq. 19, there is an additional term with $\frac{\mathbf{n}}{\mathbf{n}\cdot\mathbf{p}}$, where the denominator $\mathbf{n}\cdot\mathbf{r}$ is a scalar term. This effectively makes the change in the position $\frac{\partial\hat{D}}{\partial\mathbf{p}_{i}}$ to move along the direction of the normal $\mathbf{n}$. Consequently, the supervision on D-Normal directly affects the Gaussian position $\mathbf{p}$.
We can further deduce that the gradient-update on the Gaussian position pulls the position along the normal towards the surface, which achieves better reconstruction.
In view of the above proof, we conclude that it is better to do supervision on the D-Normal regularizer. In addition to the mathematical proof, we also visualize the positions of Gaussian centers from the optimized 3D scenes both with and without the D-Normal regularizer in Fig. 6, thereby providing experimental validation of the conclusion.
A.3 Implementation Details